# DiffUHaul: A Training-Free Method for Object Dragging in Images

### v1

@@ Abstract.
journal: TOGsubmissionid: 0ccs: Computing methodologies Machine learningccs: Computing methodologies Computer graphics
@@ 2. Related Work
Soon after the emergence of text-to-image diffusion models, a plethora of methods were offered for various image editing tasks (Meng et al., 2021; Avrahami et al., 2022, 2023b; Mokady et al., 2023; Tumanyan et al., 2023; Hertz et al., 2023; Kawar et al., 2023; Cao et al., 2023; Patashnik et al., 2023; Sheynin et al., 2023). However, most of these editing methods are spatially preserving (i.e., changing the object attributes and categories), and suffer from the editing tasks requiring spatial reasoning, such as object dragging (Brooks et al., 2023; Hertz et al., 2022). Localized text-to-image models, such as GLIGEN (Li et al., 2023) and BlobGEN (Nie et al., 2024), has the potential to solve the object dragging task, but their performance is far from satisfactory without specialized designs.
@@
A similar task is keypoint dragging, where users provide source and target keypoints in the image, and move the source keypoints to the target ones. For example, UserControllableLT (Endo, 2022), GANWarping (Wang et al., 2022) and DragGAN (Pan et al., 2023) employ StyleGAN (Karras et al., 2019, 2020, 2021) for editing generated images. But they work only on the narrow domain the GAN (Goodfellow et al., 2014) was trained on (e.g., human faces, churches (Yu et al., 2015)). DragDiffusion (Shi et al., 2023) propose a LoRA-based (Hu et al., 2021) method that finetunes a diffusion model given a test image and optimizes the latent noises at inference time. In contrast, our method is training-free.
@@ 4.1. Gated Self-Attention Entanglement
Figure 5. Gated Self-Attention Leakage. Given scene descriptions of two blobs: “a photo of a rabbit” and “a photo of a cat”, the standard BlobGEN model (first row) leaks the rabbit information also to the cat blob (the first row third column), while our masked version of the gated self-attention (second row) is able to disentangle the blobs (the second row third column). In addition, we can see that the gated self-attention (second column) behaves de facto as a cross-attention layer, as the vast majority of the attention is between the text tokens $T$ and the visual tokens $V$.
@@ 4.3. Extension for Real Images
In order to extend our method for dragging objects in real images, we first extract the blobs parameters as we explain later, then, we need to invert the image. However, we found that directly applying the DDIM inversion (Song et al., 2020) in a localized model is not able to preserve the details of the input image, even without classifier-free guidance (Ho, 2022).
@@ 5.1. Qualitative and Quantitative Comparison
We compared our method against the most relevant available object dragging baselines. Paint-By-Example (PBE) (Yang et al., 2022) and AnyDoor (Chen et al., 2023a) present a way of adding an object to an image. To use them for object dragging, we crop the object from the source image, apply the image inpaintng in the cropped region, and then add the object in the new location. Diffusion Self-Guidance (Diffusion SG) (Epstein et al., 2023) tackles the general image editing tasks via attention guidance, which can be tailored to object dragging. DragDiffusion (Shi et al., 2023) is designed for the task of keypoint-based dragging, which we can convert into object dragging by selecting multiple points on the source object. Finally, DragonDiffusion (Mou et al., 2023) and DiffEditor (Mou et al., 2024) directly tackle the problem of object dragging. For more details, please see the supplementary material.
@@
Based on this new dataset, we propose three evaluation metrics: foreground similarity, object traces and realism. Foreground similarity quantifies whether the source object indeed dragged to the target location without appearance changes. To this end, we crop a tight box area around the source blob $B_{s}$ in the source image $I_{s}$ and around the target blob $B_{d}$ in the target image $I^{\prime}$, respectively, and pass the crops to DINOv2 (Oquab et al., 2023) to measure the perceptual similarity after aligning them to a canonical position and masking the background. We strive to maximize this metric. To measure the object traces phenomenon, we crop a tight box area around the source blob $B_{s}$ in the source image $I_{s}$ and around the source blob $B_{s}$ in the target image $I^{\prime}$. Next, we masked the target blob $B_{d}$ area in the target image $I^{\prime}$. Similarly, we utilize DINOv2 (Oquab et al., 2023) to measure the perceptual similarity between the crops. We strive to minimize this metric. Lastly, to measure the realism of the edited image, we utilized KID score (Binkowski et al., 2018) of sets of 672 real and generated images. For more details, please read the supplementary material.
@@ 5.2. User Study
We conduct an extensive user study using the Amazon Mechanical Turk (AMT) platform (Amazon, 2024), where the test examples are also sampled from the automatically extracted dataset as explained in Section 5.1. We compare all the baselines using the standard two-alternative forced-choice format. Users were given the source image, the edit instructions and two edited images: one from our method and another one from a baseline. For each comparison, users were asked to rate which edited image is better in terms of: (1) dragging the object to the desired location (2) leaving no traces of the original object, (3) realism and (4) overall edit quality (i.e., taking all the aspects into account). As can be seen in Table 3, our method is preferred over all the baselines in terms of the overall edit quality and different individual perspectives. This observation aligns well with our automatic metrics. The user study suggests that DragDiffusion is the second-strongest baseline, it may be due to the fact that it also result with realistic images, as it also avoids leaving traces of the dragged object, which the automatically calculated KID do not take into account. For more details and statistical significance analysis, please read the supplementary material.
@@ 6. Limitations and Conclusions
Figure 7. Limitations. Our method suffers from the following limitations: (a) We found our method to be incapable of rotating objects, and instead stretch the object to fit the new blob shape without changing the orientation. (b) We found our method to struggle with resizing object, especially in large resizes (e.g., Resize 3 in the second row). (c) We found our method to struggle to handle colliding object while dragging, which may result with a hybrid between the objects (e.g., Drag 2 in the third row) or one object being merged (e.g., Drag 3 in the third row).
@@
In conclusion, we presented DiffUHall, our solution to the seemingly straightforward task of object dragging. We demonstrated that the spatial understanding of the localized BlobGEN can be harnessed to this task, using our novel diffusion anchoring technique that manages to merge the location signal from the model with the object appearance signal from the input image. We hope that our contributions will serve as a valuable tools not only for object dragging but also for other creative tasks in the future.
@@ A.1.1. Gated Self-Attention Visualization Implementation Details
As explained in Section 4.1, we offer a method to visualize the gated self-attention layer of GLIGEN (Li et al., 2023) and BlobGEN (Nie et al., 2024). In gated self-attention, a projection layer first converts the CLIP (Radford et al., 2021) text embeddings of the text description $S_{i}$ to the text tokens $T=\{t_{1},...t_{n}\}$. They are then merged with the visual tokens $V=\{v_{1},...v_{k}\}$ into a unified set $V\cup T=\{v_{1},...v_{k},t_{1},...t_{n}\}$, which altogether are used to calculate the self-attention features, using the standard self-attention mechanism (plus a gated skip connection).
@@ Appendix B Additional Results
Figure 14. Qualitative Results in Ablation Study. As explained in Section 5.3, we ablate four key components of our method: (a) w/o GSA masking, (b) w/o SA sharing, (c) w/o soft attention anchoring and (d) w/o DDPM SA attention. As can be seen, all these components improve the foreground object consistency.
@@
In addition, in Figure 14 we provide a qualitative visualization of the ablation study we conducted. As can be seen, PBE (Yang et al., 2022), DiffusionSG (Epstein et al., 2023) and Anydoor (Chen et al., 2023a) mainly suffer from a bad preservation of the foreground object. DragDiffusion (Shi et al., 2023) struggles with dragging the object, while DragonDiffusion (Mou et al., 2023) and DiffEditor (Mou et al., 2024) suffers from object traces. Our method, on the other hand, strikes the balance between dragging the object and preserving its identity.

### v2

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: SIGGRAPH Asia 2024 Conference Papers; December 3–6, 2024; Tokyo, Japanbooktitle: SIGGRAPH Asia 2024 Conference Papers (SA Conference Papers ’24), December 3–6, 2024, Tokyo, Japandoi: 10.1145/3680528.3687590isbn: 979-8-4007-1131-2/24/12submissionid: 308ccs: Computing methodologies Machine learningccs: Computing methodologies Computer graphics
@@ 2. Related Work
Soon after the emergence of text-to-image diffusion models, a plethora of methods were offered for various image editing tasks (Meng et al., 2021; Avrahami et al., 2022, 2023b; Mokady et al., 2023; Tumanyan et al., 2023; Hertz et al., 2023; Kawar et al., 2023; Cao et al., 2023; Patashnik et al., 2023; Sheynin et al., 2023). However, most of these editing methods are spatially preserving (i.e., changing the object attributes and categories), and suffer from the editing tasks requiring spatial reasoning, such as object dragging (Brooks et al., 2023; Hertz et al., 2022). Localized text-to-image models, such as GLIGEN (Li et al., 2023) and BlobGEN (Nie et al., 2024), has the potential to solve the object dragging task, but their performance is far from satisfactory without specialized designs. Concurrently, Diffusion Handles (Pandey et al., 2023) offers 3D object edits using a depth-to-image diffusion model and by performing manipulations on the diffusion activations in 3D. In addition, Magic Fixup (Alzayer et al., 2024) offers a model that given a coarsely edited image, synthesizes a photorealistic version of it, by leveraging a video dataset, this way they manage to offer a way to edit an image coarsely, and then harmonize the result.
@@
A similar task is keypoint dragging, where users provide source and target keypoints in the image, and move the source keypoints to the target ones. For example, UserControllableLT (Endo, 2022), GANWarping (Wang et al., 2022) and DragGAN (Pan et al., 2023) employ StyleGAN (Karras et al., 2019, 2020, 2021) for editing generated images. But they work only on the narrow domain the GAN (Goodfellow et al., 2014) was trained on (e.g., human faces, churches (Yu et al., 2015)). DragDiffusion (Shi et al., 2023) propose a LoRA-based (Hu et al., 2021) method that finetunes a diffusion model given a test image and optimizes the latent noises at inference time. In contrast, our method is training-free. Concurrently, EasyDrag (Hou et al., 2024) improves DragDiffusion (Shi et al., 2023) by replacing the LoRA training with reference guidance.
@@ 4.1. Gated Self-Attention Entanglement
Figure 5. Gated Self-Attention Leakage. Given scene descriptions of two blobs: “a photo of a rabbit” and “a photo of a cat”, we can see that the standard BlobGEN model (the first column in the first row) generates two rabbits instead of a cat and a rabbit, we then visualize the gated self-attention layers, as explained in Section 4.1. As can be seen, the standard BlobGEN model (first row) leaks the rabbit information also to the cat blob (the first row third column), while our masked version of the gated self-attention (second row) is able to disentangle the blobs (the second row third column). In addition, we can see that the gated self-attention (second column) behaves de facto as a cross-attention layer, as the vast majority of the attention is between the text tokens $T$ and the visual tokens $V$.
@@ 4.3. Extension for Real Images
In order to extend our method for dragging objects in real images, we first extract the blobs parameters as we explain later, then, we need to invert the image. However, we found that directly applying theg DDIM inversion (Song et al., 2020) in a localized model is not able to preserve the details of the input image, even without classifier-free guidance (Ho, 2022). Using more advanced inversion methods (Huberman-Spiegelglas et al., 2023; Qi et al., 2023) will not work in our case, as they preserve the general structure of the scene, whereas we are interested in changing the scene layout significantly.
@@ 5.1. Qualitative and Quantitative Comparison
We compared our method against the most relevant available object dragging baselines. Paint-By-Example (PBE) (Yang et al., 2022) and AnyDoor (Chen et al., 2023a) present a way of adding an object to an image. To use them for object dragging, we crop the object from the source image, apply the image inpainting in the cropped region, and then add the object in the new location. Diffusion Self-Guidance (Diffusion SG) (Epstein et al., 2023) tackles the general image editing tasks via attention guidance, which can be tailored to object dragging. DragDiffusion (Shi et al., 2023) is designed for the task of keypoint-based dragging, which we can convert into object dragging by selecting multiple points on the source object. Finally, DragonDiffusion (Mou et al., 2023) and DiffEditor (Mou et al., 2024) directly tackle the problem of object dragging. For more details, please see the supplementary material.
@@
Based on this new dataset, we propose three evaluation metrics: foreground similarity, object traces and realism. Foreground similarity quantifies whether the source object indeed dragged to the target location without appearance changes. To this end, we crop a tight box area around the source blob $B_{s}$ in the source image $I_{s}$ and around the target blob $B_{d}$ in the target image $I^{\prime}$, respectively, and pass the crops to DINOv2 (Oquab et al., 2023) to measure the perceptual similarity after aligning them to a canonical position and masking the background. We strive to maximize this metric. To measure the object traces phenomenon, we crop a tight box area around the source blob $B_{s}$ in the source image $I_{s}$ and around the source blob $B_{s}$ in the target image $I^{\prime}$. Next, we mask the target blob $B_{d}$ area in the target image $I^{\prime}$. Similarly, we utilize DINOv2 (Oquab et al., 2023) to measure the perceptual similarity between the crops. We strive to minimize this metric. Lastly, to measure the realism of the edited image, we utilize KID score (Binkowski et al., 2018) of sets of 672 real and generated images. For more details, please read the supplementary material.
@@ 5.2. User Study
We conduct an extensive user study using the Amazon Mechanical Turk (AMT) platform (Amazon, 2024), where the test examples are also sampled from the automatically extracted dataset as explained in Section 5.1. We compare all the baselines using the standard two-alternative forced-choice format. Users were given the source image, the edit instructions and two edited images: one from our method and another one from a baseline. For each comparison, users were asked to rate which edited image is better in terms of: (1) dragging the object to the desired location (2) leaving no traces of the original object, (3) realism and (4) overall edit quality (i.e., taking all the aspects into account). As can be seen in Table 3, our method is preferred over all the baselines in terms of the overall edit quality and different individual perspectives. This observation aligns well with our automatic metrics. The user study suggests that DragDiffusion is the second-strongest baseline, it may be due to the fact that it also results with realistic images, as it also avoids leaving traces of the dragged object, which the automatically calculated KID do not take into account. For more details and statistical significance analysis, please read the supplementary material.
@@ 6. Limitations and Conclusions
Figure 7. Limitations. Our method suffers from the following limitations: (a) We found our method to be incapable of rotating objects, and instead stretch the objects to fit the new blob shape without changing the orientation. (b) We found our method to struggle with resizing objects, especially in large resizes (e.g., Resize 3 in the second row). (c) We found our method to struggle to handle colliding objects while dragging, which may result with a hybrid between the objects (e.g., Drag 2 in the third row) or one object being merged (e.g., Drag 3 in the third row).
@@
In conclusion, we presented DiffUHaul, our solution to the seemingly straightforward task of object dragging. We demonstrated that the spatial understanding of the localized BlobGEN can be harnessed to this task, using our novel diffusion anchoring technique that manages to merge the location signal from the model with the object appearance signal from the input image.
Acknowledgements.
This work was supported in part by the Israel Science Foundation (grants 2492/20, 3611/21 and 1574/21).
@@ A.1.1. Gated Self-Attention Visualization Implementation Details
As explained in Section 4.1 we offer a method to visualize the gated self-attention layer of GLIGEN (Li et al., 2023) and BlobGEN (Nie et al., 2024). In gated self-attention, a projection layer first converts the CLIP (Radford et al., 2021) text embeddings of the text description $S_{i}$ to the text tokens $T=\{t_{1},...t_{n}\}$. They are then merged with the visual tokens $V=\{v_{1},...v_{k}\}$ into a unified set $V\cup T=\{v_{1},...v_{k},t_{1},...t_{n}\}$, which altogether are used to calculate the self-attention features, using the standard self-attention mechanism (plus a gated skip connection).
@@ Appendix B Additional Results
Figure 14. Qualitative Results in Ablation Study. As explained in Section 5.3, we ablate four key components of our method: (a) w/o GSA masking, (b) w/o SA sharing, (c) w/o soft attention anchoring and (d) w/o DDPM SA attention. As can be seen, all these components improve the foreground object consistency. For example, see the distorted face of the zebra or the distorted shape of the water pipe in the ablated cases.
@@
In addition, in Figure 14 we provide a qualitative visualization of the ablation study we conducted. As can be seen, all these components improve the foreground object consistency. Please notice, the contribution of the GSA masking, as reflected by our automatic metric in Table 2 in the main paper, is limited in comparison to the other components that we added. However, the GSA masking mitigates the GSA leakage, which affects changes in the identity of the moved objects. This effect is visualized in Figure 14: see the distorted face of the zebra or the distorted shape of the water pipe in the second row, which ablates GSA masking.